The center of a circular sunflower with a diameter of 4 centimeters is (-2,1). Write the equation that represents the sunflower.
Please help & Thanks!

Answer by scott8148(6628)

(x-(-2))^2+(y-1)^2=(4/2)^2 __ x^2+4x+4+y^2-2y+1=4 __ x^2+y^2+4x-2y+1=0